# A Dataset and Benchmark for Copyright Protection from Text-to-Image Diffusion Models

### v1

@@ 5 CPDM Benchmark
5.1 Gradient Ascent-based Approach
To make diffusion models forget a specific copyrighted image, a simple and effective method is to train the model using gradient ascent optimization on that image. For a single image, forgetting can be achieved by optimizing for a few epochs with an appropriate learning rate. More specifically, for a diffusion model with its set of weight parameters $\theta$, to forget the image $Y$ and its corresponding prompt $X$, we update $\theta$ each epoch in the following way:
$\theta=\theta+\eta\nabla_{Y}L_{mse}(\theta,X,Y)$
@@ 5.1 Gradient Ascent-based Approach
5.2 Weight Pruning-based Approach
Weight pruning [8, 17, 18, 28] is an effective method for reducing a model’s parameter count, commonly utilized in model deployment and practical applications. Moreover, this method can be adapted to modify model parameters, enabling the model to forget specific copyrighted images. Inspired by magnitude pruning [17, 18], the core idea behind parameter pruning for forgetting certain infringing images is to mask the weights in the model so that those weights exhibit the strongest response in generating those images. In our experiments, we first feed the image to be forgotten into Stable Diffusion for forward propagation, simultaneously obtaining the gradients of each layer in the network. For the pruning stage, we regard each layer as an individual pruning group. The highest $p_{c}\%$ of activation values are identified within each layer, and we locate the weights correlated with these values. Subsequently, based on the magnitudes of the gradients of these weights, we set the top $p_{w}\%$ of weights to zero. This process can be described by the following equation:
@@ 5.2 Weight Pruning-based Approach
To illustrate, for a layer expressed as $Y=WX$ where $W$ represents the weight (bias term omitted for simplicity), we first select $W_{i}$ corresponding to the greatest $p_{c}\%$ of $|Y_{ij}|$, where ||$\cdot$|| ⋅ | represents the absolute value operator, $optim\{\cdot\}$ represents updating parameters. Then, for each $W_{i}$, we prune $p_{w}\%$ of the elements $W_{ij}$ corresponding to the highest $m\%$ of its gradient values $\nabla_{W}L_{ij}$, setting these to zero.
6 Experiments
Table 3: Benchmark Performance Testing on data from WikiArt and Wikipedia. The CPDM metrics are evaluated on a case-by-case basis, comparing anchored images to the generated images produced by respective models. For images generated by SD-v2.1, we amalgamated copyright-attribute images with the images generated by SD-v2.1, forming an “Image Pair”. In the context of the unlearning process, we paired copyright images with images generated after applying our proposed “Prune” and “Gradient Ascent” methods, as well as with those produced using “ESD”[9] and “Forget-Me-Not”[42] methods for comparative analysis (further details on the setups for ESD and Forget-Me-Not are provided in the supplementary materials). To elaborate, we calculated the “FID” of the generated images from SD-v2.1 and its unlearned counterparts using the COCO-10K dataset [27]. Furthermore, we uniformly assessed the “CM” (an abbreviation for our proposed “CPDM” metric for distinction) and $\Delta$CLIP” metrics to elucidate their effectiveness. CPDM Metric demonstrates a robust measure of the ability to forget copyright-protected images.
SD-Generated Prune Gradient Ascent ESD Forget-Me-Not
[…]
CM(%) $\downarrow$ 98.14 46.80 54.01 95.79 71.93
$\Delta$CLIP(%) $\downarrow$ / -21.93 -19.77 -36.24 -41.90
FID $\downarrow$ 11.18 11.34 11.79 15.72 12.43
In the Experiments, we systematically analyze the effectiveness of the proposed CPDM metrics, and provide the results and analysis of benchmark methods on our proposed CPDM datasets.
6.1 Experimental Setting
We conducted experiments on our proposed dataset using two baseline methods. Specifically, we performed image forgetting experiments on the Wikiart, Portrait, Cartoon, and Illustration domains. We calculated the corresponding metrics to evaluate the effectiveness of image forgetting and used the FID (Fréchet Inception Distance) [7, 19] to assess the change in the generative capability of the models after unlearning. In the following two unlearning algorithms, we only made parameter adjustments for the UNet structure of Stable Diffusion [33]. We froze the parameters of the text embedding module and the autoencoder module. The respective experiments were conducted on stable diffusion (SD-v1.4), finetuned diffusion model (SD-finetuned), and (SD-v2.1). We performed unlearning experiments on each image in the CPDM dataset. For calculating FID for each image category, we randomly selected five post-unlearning models. Subsequently, we generated 10,000 images using the diffusion model and used them to compute FID in comparison with the COCO-10k dataset [27].
Table 4: Benchmark Performance Testing on data from anonymous artist, illustration (Illu.) part of the dataset. The “SD-Generated” and “SD-finetuned” stand for the SD-V1.4 and its finetuned counterpart on our Illustraion dataset.
SD-Generated SD-finetuned Prune Gradient Ascent ESD Forget-Me-Not
[…]
CM(%) $\downarrow$ 97.92 99.69 95.74 94.57 95.90 83.24
$\Delta$CLIP(%) $\downarrow$ 5.08 / -21.56 -31.18 -2.08 -16.17
FID $\downarrow$ 13.21 9.33 36.93 11.85 8.61 12.95
6.2 Metrics Effectiveness Evaluation
Figure 4: Visualization results of our metrics for 100 selected generated images and corresponding anchor images: (a) represents the Semantic Metric. (b) represents the Style Metric, and (c) represents the Total CPDM Metric. Brighter pixels indicate higher similarity between the two images, as determined by the formula $Softmax(-M^{100\times 100})$. (c) demonstrates that the highlighted pixels are predominantly distributed near the diagonal line, effectively validating that our CPDM metric can comprehensively assess the similarity between two images. In addition, as ablation studies, (a) and (b) reveal that the Semantic metric is not sensitive to certain dissimilarities, while the Style metric is overly sensitive to some dissimilarities. The $w^{1~{}n}$ we choose is $[0.5,0.1,6e4,4]$
To validate our metric’s effectiveness, we first visualize the results it produces. In this experiment, we selected 10 artists and generated images that mimic the style of their artworks, using the pipeline introduced in Section 4. We then randomly selected 10 generated images per artist, with visually similar features to their anchor image (potentially indicative of plagiarism), in total of 100 images. With these selections, we utilize our metrics to compute a matrix $M^{100*100}$, capturing the relationships between the 100 generated images and their corresponding 100 anchor images. The distributions of Semantic Metric, Style Metric and Total CPDM Metric are visualized in the Fig. 4. These visualizations evidently support the assertion that our metric can successfully identify images that may constitute plagiarism. Furthermore, it can also apply for the metrics in unlearning task, which allows for a more systematic and quantifiable assessment of the methods for unlearning.
@@ 6.2 Metrics Effectiveness Evaluation
Figure 5: Experimental results of unlearning algorithm on specially finetuned generation models.
Figure 6: The copyright-infringing images generated by the generative model. Top row: Original copyrighted image. Middle row: Output from SD-finetuned. Bottom row: Output from SD-v2.1.
6.3 Results and Analysis
Benchmark: Gradient Ascent-based Approach
We have two primary foundational models in our repertoire. The first one is our finetuned model, SD-finetuned, which is employed to evaluate and assess the performance of the unlearning algorithm on the finetuned full parameters of the UNet within the stable diffusion framework. This evaluation includes both the model’s ability to forget copyrighted images and its performance in generating other types of images. The second foundational model is an unlearning experiment conducted on the stable diffusion v2.1, which is currently the best publicly available generative model based on stable diffusion. During the unlearning experiments conducted on these two foundational models, for each image to be forgotten, the learning rate for gradient ascent is set at 5.0e-05, and unlearning training is performed for five epochs.
Benchmark: Weight Pruning-based Approach
Similar to gradient ascent, we conducted unlearning experiments on both foundational models using our dataset. For the SD-finetuned model, which has been fine-tuned on specific illustration styles, we set the pruning ratios $p_{c}\%$ and $p_{w}\%$ to 0.1 and 0.03, respectively. As for the SD-v2.1 model, we set the pruning ratios $p_{c}\%$ and $p_{w}\%$ to 0.1 and 0.005, respectively. The reason for employing a higher pruning ratio on SD-finetuned is due to the need for a more significant pruning ratio to forget such artistic styles when the model has been fine-tuned on a limited amount of data. For each image to be forgotten, we performed one epoch of iterative computation and pruning. During the pruning process, the optimizer remained disabled.
Infringement Model: SD-Finetuned By combining the prompts obtained from the pipeline described in Fig.3, we can simulate a model that utilizes stable diffusion to infringe upon specific artworks. Throughout the training phase of the SD-finetuned model, intended for generating infringing illustrations, we meticulously employed a set of 160 artwork images in conjunction with their corresponding prompts. To ensure a more faithful simulation of the process of infringing upon artwork, we opted for a two-stage fine-tuning procedure. This meticulous approach was implemented to yield infringing images that bear a striking resemblance to the original artworks in terms of their artistic characteristics, encompassing style, lines, lighting, composition, brushstrokes, and more. We shall meticulously refine the complete spectrum of parameters within the Unet model to replicate the intricate nuances of infringement within the realm of artwork. The specific training details are provided in the appendix.

### v3

@@ 3.1 Dataset Creation Pipeline
Figure 4: Analysis of Dataset Style Diversity. We conducted a statistical of the styles in prompts obtained using the pipeline in Fig. 3. The various styles at the bottom are provided by artists.
3.2 Dataset Composition
After engaging in thorough discussions with both artists and legal experts, we divided the dataset into the following four parts, to facilitate a comprehensive examination from various viewpoints to identify and address potential copyright infringement concerns effectively. The sources of our curated data are listed in Tab. 1. The style diversity statistics of the dataset are shown in Fig. 4.
Style. Painting artworks often embody the distinctive style of the artist, encompassing aspects such as brushstrokes, lines, colors, and compositions. This artistic style is also a form of copyright that requires protection. WikiArt is an online user-editable visual art encyclopedia, a source from which numerous art-related datasets [24, 29] have been curated. WikiArt already features some 250,000 artworks by 3,000 artists, localized on 8 languages. These artworks are in museums, universities, town halls, and other civic buildings of more than 100 countries. We selected approximately 1500 artworks from 100 artists on WikiArt as the source of anchor images for prompt generation and corresponding content generation using stable diffusion.
Portrait. The right of portrait refers to an individual’s control and use of their own portrait, including facial features, image, and posture. The purpose of publicity rights is to safeguard an individual’s privacy, personal dignity, and image integrity, preventing unauthorized use, disclosure, or alteration of their portrait. This legal protection aims to preserve an individual’s control over how their likeness is portrayed and commercialized. We utilized web scraping techniques to collect over 200 portrait images from Wikipedia, which is a free, web-based, multilingual encyclopedia that contains articles on a wide range of topics.
Artistic Creation Figure. Artistic creations, including characters from animations and cartoons, are often protected by law. In this context, we refer to this category as “artistic creation figures”. Similar to portraits, we have curated a dataset of 200 influential animated characters and figures by collecting information from reputable sources such as Wikipedia.
Licensed Illustration. We have obtained authorization to use a portion of Anonymous Artist’s artworks in this study. Therefore, we can include these artworks as part of the training dataset for fine-tuning stable diffusion, which will be utilized for simulating infringing artistic paintings.
@@ 4 CPDM Benchmark
4.1 CPDM Metric
Table 2: Comprehensive Evaluation Metrics. In this context, “Imagecpr, “Promptcpr, “Modelori, “Modelunl, and “Modelicp represent the copyright image, the prompt associated with the copyright image, the original SD model, the SD model after unlearning, and the Inception model, respectively.
Input Metric Description
Imagecpr, Promptcpr Imagegen: Modelori (Promptcpr) Imageunl: Modelunl (Promptcpr) Modelori, Modelunl, Modelicp Potential for Infringement CM <Imagecpr, Imagegen >
Effectiveness of Unlearning CM <Imagecpr, Imageunl >
$\Delta$CLIP
<Imagegen, Promptcpr>
<Imageunl, Promptcpr>
Model Degradation FID Modelicp (Imageunl)
In Tab. 2, the metric consists of CM, $\Delta$CLIP, and FID. As the latter two metrics are apparent in formulation, we focus on illustrating the construction of the CM metric in this section. We collaborate with an artist, Anonymous Artist, who is currently active in the art industry and is notably distinguished in the field of illustration. This artist offers invaluable insights from a professional standpoint to help assess whether two images constitute plagiarism, considering elements such as brushstrokes, color palettes, lighting effects, and composition. However, accurately defining these potentially copyrighted attributes of images is challenging. We initially divide measurements into two perspectives: semantic and stylistic components. The goal is to develop a formula that combines these two components and provides a scalar measure to quantify the similarity between two images.
Semantic Metric. We leverage the CLIP[32] model to generate the semantic embedding, and calculate the metrics by:
$emb_{ori}=E_{clip}(Image_{ori}),\quad emb_{gen}=E_{clip}(Image_{gen})$ (1)
$M_{\text{sem}}=\frac{1}{n}\sum_{i=1}^{n}(emb_{\text{ori}}[i]-emb_{\text{gen}}[i])^{2}$ (2)
where $Image_{ori}$ and $Image_{gen}$ denote the anchor image and generated image respectively; $E_{clip}$ denotes the CLIP’s image encoder; $i$ represents pixel value. In previous studies, cosine similarity has been predominantly used as the evaluation metric. However, in this research, we utilize MSE instead. This choice is motivated by two primary factors: first, the range of MSE is significantly broader than that of cosine similarity, making it easier to observe changes resulting from unlearning; second, adopting MSE aligns better with the subsequent style metrics discussed below.
Style Metric. Inspired by [10], we use the activation output of CNN networks to calculate feature correlations via the Gram matrix. We leverage InceptionV3 [39], following the FID metric.
$G^{l}_{ori}=Gram(E_{incep}(Image_{ori},l)),\quad G^{l}_{gen}=Gram(E_{incep}(Image_{gen},l))$ (3)
$D^{l}=\frac{1}{n}\sum_{i=1}^{n}(G^{l}_{ori}[i]-G^{l}_{gen}[i])^{2}$ (4)
$M_{style}=\sum_{i=1}^{n}w^{l}D^{l}$ (5)
where $E_{incep}(Image,\,l)$ denotes passsing the $Image$ through an Inception network and extracting the feature maps from layer $l$. The Gram matrix is then computed to provide a style representation of the image at layer $l$. Furthermore, the dissimilarity between the original and generated images in each layer is represented by MSE of the Gram matrices in each corresponding layer. The total style metric, as described above, is determined by weighting factors $w^{l}$, which represent the contribution of each layer to the overall style metrics. In our work, $n$ is set to $4$, because there are four stages of the InceptionV3 model. The values of parameters $w^{l}$ needs to be fine-tuned according to the distribution of the images. We provide specific values applicable to our dataset in the supplementary material. Finally, we denote the total metric as:
$CM$ $=1-Norm(M_{sem}\times M_{style})^{2}$ (6)
Here, $Norm$ represents normalization, which involves clipping data to the range $[a,b]$, then dividing by $(b-a)$, with $a$ and $b$ empirically set to 1 and 50, respectively. The squared term emphasizes the significant changes before and after the unlearning process. The effectiveness of this quantifiable metric will be verified in section 5.2.
4.2 Benchmark Unlearning Methods
To facilitate fair comparison among existing unlearning methods using our proposed CPDM benchmark, we present two intuitive baseline unlearning methods: a gradient ascent-based approach and a weight pruning-based approach.
Gradient Ascent-based Approach (GA). For a single image, forgetting can be achieved by optimizing for a few epochs with an appropriate learning rate [13, 23]. More specifically, for a diffusion model with its set of weight parameters $\theta$, to forget the image $Y$ and its corresponding prompt $X$, we update $\theta$ each epoch in the following way:
$\theta=\theta+\eta\nabla_{\theta}L_{mse}(\theta,X,Y)$
@@ 4.2 Benchmark Unlearning Methods
Weight Pruning-based Approach (WP). Inspired by magnitude pruning [17, 18], the core idea behind parameter pruning for forgetting certain infringing images is to mask the weights in the model so that those weights exhibit the strongest response in generating those images. Initially, the image is passed through Stable Diffusion for forward propagation, simultaneously obtaining gradients for each network layer. Each layer is treated as a separate pruning group. The highest $p_{c}\%$ activation values within each layer are identified, and weights correlated with these values are located. Subsequently, based on the gradient magnitudes of these weights, the top $p_{w}\%$ are set to zero. This process can be described by the following equation:
@@
To illustrate, for a layer expressed as $Y=WX$ where $W$ represents the weight (bias term omitted for simplicity), we first select $W_{i}$ corresponding to the greatest $p_{c}\%$ of $|Y_{ij}|$, where ||$\cdot$|| ⋅ | represents the absolute value operator, $optim\{\cdot\}$ represents updating parameters. Then, for each $W_{i}$, we prune $p_{w}\%$ of the elements $W_{ij}$ corresponding to the highest $m\%$ of its gradient values $\nabla_{W}L_{ij}$, setting these to zero. More related research work and experiments can be found in the supplementary materials.
5 CPDM Experiments
5.1 Experiment Setting
Our experiments can be roughly categorized into two perspectives. First, indicating the potential of infringement of SD-Generated images, using the original version of SD (SD-v1.4, SD-v2.1) and the finetuned counterpart (SD-Finetuned) respectively. Note that we directly test the SD generative models using the data from WikiArt and Wikipedia, as shown in Tab. 3’ top. We test both the SD-Generated and SD-Finetuned models on the Illustration part of our CPDM dataset, from an anonymous artist, as shown in Tab. 3’ bottom. Second, we conduct a comprehensive evaluation of unlearning methods, regarding their ability to remove copyright-relevant concepts and maintain the copyright-irrelevant generative capacities. In this stage, we incorporate four representative unlearning methods tailored for text-to-image generative tasks, ESD[7], FMN[42], CA[27], and UCE[8] (further details on the setups are provided in the supplementary materials), along with our proposed baseline unlearning methods, GA and WP. As shown in Tab 3, the CPDM metrics are evaluated on a case-by-case basis, comparing anchored copyright images to the generated images produced by SD models and their unlearned counterparts, forming image pairs of <Imagecpr, Imagegen> and <Imagecpr, Imageunl>, respectively. Specifically, FID is calculated by generating 10,000 images using each model on the COCO-10K dataset [30]. In the context of an unlearning process, for our proposed two baseline unlearning methods, we only make parameter adjustments on the U-Net structure of SD models[34], while freezing the other parameters.
5.2 Benchmark Metric Evaluation
Figure 5: Analysis and Comparison of the Effectiveness of Evaluation Metrics.
In Fig. 5, we selected ten artists and generated images that mimic their artwork styles using the pipeline described in Fig. 3. For each artist, ten generated images were randomly chosen, resulting in a total of 100 images. We then used our metrics to compute a matrix ($M^{100\times 100}$), capturing the relationships between the generated images and their corresponding anchor images. The horizontal axis representing anchor images and the vertical axis representing generated images. Brighter pixels indicate higher similarity between image pairs. A brighter diagonal suggests a more accurate metric assessment. In (c), the highlighted pixels primarily near the diagonal confirm the CM metric’s effectiveness in assessing image similarity. Additionally, (a) and (b) illustrate the Semantic metric’s insensitivity to certain dissimilarities, while the Style metric is overly sensitive to some variations. The results of (c) and (d) also demonstrate the rationality of the design of the style metric in Section 4.1. These visualizations support the assertion that our metric can identify images that may constitute plagiarism. Moreover, this approach can be applied to unlearning tasks, providing a systematic and quantifiable method for evaluating unlearning techniques.
Figure 6: Consistency analysis between human evaluation and CM metric.
5.3 Human Verification
To ensure a high-quality dataset and the design of a more compelling evaluation metric, we manually verified the dataset obtained from the creation pipeline and assessed the consistency between CM and manual evaluation. 1) For the CPDM dataset, as the current dataset is relatively small and manageable for manual screening, we assembled a team of four individuals with expertise in image-related work to verify the dataset. These experts meticulously reviewed the dataset, removing dissimilar images. To ensure a high level of similarity among the remaining images, a final vote was conducted to decide the inclusion of each image in the dataset. 2) For the CM metric, as shown in Fig. 6, we analyzed 100 sets of images with 119 manual evaluation results. The details of the experimental setup can be seen in the gray table in the image. The experimental results indicate that our metrics achieve results fairly consistent with manual evaluations in determining image similarity.
Table 3: Top: benchmark performance on WikiArt and Wikipedia. Bottom: benchmark performance on CPDM Illustration. On the right, WP and others represent different unlearning methods.
Metric SD-Generated SD-Finetuned WP GA ESD FMN CA UCE
<Imagecpr, Imagegen > <Imagecpr, Imageunl >
Top CM(%) $\downarrow$ 98.14 / 46.80 54.01 85.79 71.93 97.11 87.45
$\Delta$CLIP(%) $\downarrow$ / / -21.93 -19.77 -36.24 -41.90 -9.67 -20.33
FID $\downarrow$ 11.18 / 11.34 11.79 15.72 12.43 20.01 17.50
Bottom CM(%) $\downarrow$ 91.92 99.69 95.74 94.57 85.90 83.24 94.52 92.57
$\Delta$CLIP(%) $\downarrow$ 5.08 / -21.56 -31.18 -2.08 -16.17 -9.76 -0.63
FID $\downarrow$ 13.21 9.33 36.93 11.85 8.61 12.95 25.35 27.43
@@ 5.3 Human Verification
Figure 7: Experimental results of unlearning method on specially finetuned diffusion models.
5.4 Analysis and Interpretation
Interpretation. As shown in Fig. 7, visualizations of the experimental results for various unlearning algorithms are provided, along with the infringement images generated by the SD-Finetuned. 1) GA. This experiment utilized two foundational models, SD-Finetuned and the stable diffusion v2.1. They were respectively employed to evaluate the effectiveness of unlearning algorithms on customized fine-tuned generative models and basic open-source models. During the unlearning experiments, for each image to be forgotten, the learning rate for GA is set at 5.0e-05, and unlearning training is performed for five epochs. 2)WP. For the SD-Finetuned model, which is based on SD-v1.4 and has been fine-tuned on specific illustration styles, we set the pruning ratios $p_{c}\%$ and $p_{w}\%$ to 0.1 and 0.03, respectively. As for SD-v2.1 models, we set the pruning ratios $p_{c}\%$ and $p_{w}\%$ to 0.1 and 0.005, respectively. The reason for employing a higher pruning ratio on SD-Finetuned is due to the need for a more significant pruning ratio to forget such artistic styles when the model has been fine-tuned on a limited amount of data. For each image to be forgotten, we performed one epoch of iterative computation and pruning. During the pruning process, the optimizer remained disabled. 3) SD-Finetuned. During the fine-tuning process of the SD-Finetuned model, we carefully utilized a set of 160 art images along with their corresponding prompts. To more accurately simulate the process of generating infringing artworks, we opted for a two-stage fine-tuning procedure. Implementing this meticulous approach allows us to generate infringing images that exhibit high similarity to the original artworks in artistic features and various image styles. Specific training details are provided in the appendix.
Analysis. The bottom of Tab.3 represents the results on the SD-Finetuned model using CPDM Illustration data, corresponding to the visualizations in Fig.7. A detailed analysis of this part of the experiment reveals the following: 1) The CM metric for images generated directly by the diffusion model is 91.92, while for images generated by SD-Finetuned, the CM metric is as high as 99.69. This indicates that general generative models cannot directly replicate pictures of a specific style, but after fine-tuning, they can imitate the image style very closely. 2) The unlearning effects of the baseline methods WP and GA are relatively weak. As shown in Fig.7, the results of WP and GA still contain illustration styles similar to the copyrighted images. Additionally, the WP method significantly affects the model’s generative capability (refer to the FID changes). 3) Comparing several public unlearning methods, ESD, FMN, CA, and UCE, it is found that FMN has a better forgetting effect (lower CM in Tab.3). Furthermore, it can be observed that the images generated by FMN in Fig.7 differ significantly in artistic style from the copyrighted images. 4) The forgetting effect of CA is weaker compared to other methods (higher CM metric in Tab.3). Observing the images generated by CA in Fig.7, CA’s images feature a representative circular illustration pattern surrounding a perfume bottle, similar to the copyrighted images. The evaluation metrics and visualization images illustrate the effectiveness of the benchmark.